# Unleashing the Potential of Large Language Models for Text-to-Image Generation through Autoregressive Representation Alignment

### v2

@@ 3.3 Autoregressive Representation Alignment
We align the visual representations extracted from the pretrained foundational model with LLM’s hidden states and investigate impact of different alignment strategies in Section 4.2. The goal of alignment is to enable the hidden state of the autoregressive transformer to acquire external global feature representations, providing meaningful guidance for reconstructing image.
Pre-trained Visual Representation Extraction. Formally, let $\mathcal{E}_{F}$ be a pretrained foundation model’s encoder and consider a target image $I$. We encode $I$ as a visual representation by $f_{F}=\mathcal{E}_{F}(I)\in\mathbb{R}^{N\times D}$, where $N,D$ denotes the embedding length and dimension of $f_{F}$. Next, $f_{F}$ is aggregated to the global visual representation $f_{GF}$, i.e., $f_{GF}=\text{agg}(f_{F})\in\mathbb{R}^{1\times D}$, where $\text{agg}(\cdot)$ denotes a feature aggregation operation. This aggregation operation fully extracts the global information in the features and facilitates alignment with the feature of autoregressive model. For CLIP series, inspired by [40], we use the <CLS> token representation from the Transformer-based visual encoder as global visual representation. For SAM series, which lack a <CLS> token, we instead apply average pooling over all patch features for feature aggregation operation.
Hybrid Next Token. We define the next token predicted by LLM sequence in our framework as “HYBRID next token”, denoted as <HYBNEXT>. Unlike a “locally constrained token” in previous autoregressive models that are solely constrained by the LLM codebook, our defined <HYBNEXT> can fully incorporate external, well-trained global visual representations, making it a “globally and locally constrained token.”
Global Visual Representation Alignment. Specifically, we obtain the hidden state $f_{L}^{i}$ of <HYBNEXT> from the autoregressive model $\mathcal{M}_{\theta}$. The hidden state $f_{L}^{i}$ is converted to $f_{A}\in\mathbb{R}^{1\times D}$ by a projection layer $\mathcal{A}_{\phi}$ to align with the global visual representation $f_{GF}\in\mathbb{R}^{1\times D}$, i.e., $f_{A}=\mathcal{A}_{\phi}(f_{L}^{i})$, where $\mathcal{A}_{\phi}$ is a two-layer MLP. Representation alignment is achieved through a Global Visual Alignment loss $\mathcal{L}_{\text{GVA}}$, which maximizes the similarity between the projected feature $f_{A}$ and the global visual representation $f_{GF}$:
@@
where sim($\cdot$,$\cdot$ , ⋅) denotes cosine similarity loss. This alignment enables the <HYBNEXT> token to learn global visual representation, bridging the cross-modal gap and making the token prediction process more reliable.
@@
4 Experiments and Results
We conducted extensive experiments to rigorously evaluate the performance of ARRA and analyze the impact of its components. Specifically, we answer the following critical questions to the ARRA framework:
  • Visual encoder selection on alignment: How does the choice of external visual encoder (general vs. specialized and cross-modal vs. vision-only) impact generation performance? (section 4.2.1)
  • Alignment mechanism: Does token-level alignment (our proposed [HYBNEXT]) outperform fixed-position alignment ([REP]) in preserving global visual constraints during autoregressive generation? (section 4.2.2)
  • Alignment depth: How does alignment depth (i.e., the layer in the MLLM chosen for alignment with the visual encoder) affect the final performance? (section 4.2.3)
  • Feature aggregation strategy: How do different types of features extracted from the same visual encoder affect the generation performance? (section 4.2.4)
  • Main results: Does the ARRA framework exhibit better generation performance and convergence speed compared to other generation methods? (section 4.3)
@@ 4.1 Experimental Setup
Datasets. We train and evaluate our model with MIMIC-CXR [20], DeepEyeNet [19] and ImageNet [10] datasets. For the MIMIC-CXR, we employ three distinct radiographic views: posteroanterior (PA), anteroposterior (AP), and lateral (LATENT). Input prompts are generated from the “impression” section of diagnostic reports, formatted as: “{view} view chest X-ray image, {impression}”. We select 221,238 pairs for training and 1,000 pairs for testing. DeepEyeNet [19] is a comprehensive fundus dataset, containing more than 15k high quality “text-image” pairs. In our experiments, we derive prompts from the “clinical-description” section and selectively chose 7,190 training pairs and 1,089 testing pairs. All images are preprocessed through center cropping to 512×512 pixels. ImageNet [10] is a popular natural image datasets containing more than 1.2 million samples. We randomly select 256,233 images from ImageNet-1k as our training data, and all samples were generated and evaluated in 256 × 256 resolution.
Evaluation Metrics. We employ three quantitative metrics to evaluate the proposed method: Fréchet Inception Distance (FID), Multi-Scale Structural Similarity (MS-SSIM), and CLIP-Score. For X-ray image generation, FID is computed using features extracted from a CXR-pre-trained DenseNet-121 (XRV) [8], while for fundus images, we utilize features from an ImageNet-pre-trained Inception V3 [17]. The CLIP-Score metric is derived from the cosine similarity between image-text feature pairs encoded through BioMedCLIP [61], a large-scale medical vision-language model.
Implementation Details. We validate the versatility of ARRA under three distinct configurations to demonstrate its adaptability across architectures, domains, and training paradigms. First, ARRA-Base trains from random initialization, testing ARRA’s ability to bootstrap multimodal alignment without pretrained weights. Second, ARRA initializes with a pretrained LLM that has strong text generation capabilities, evaluating its capacity to repurpose off-the-shelf LLMs for text-to-image generation. Third, ARRA-Adapt starts from a pretrained LLM with both text and image generation capabilities, testing ARRA’s ability to repurpose general-domain image generation priors for specialized tasks (e.g., medical imaging).
For MIMIC-CXR and DeepEyeNet datasets, we employ both ARRA using Chameleon without image generation capability[50] as a base LLM to train the visual generation model, and ARRA-Adapt by fine-tuning the Lumina-mGPT [27]) model with image generation capability. For ImageNet dataset, we test the ARRA-Base configuration using LLamaGen network [48]. All constructed models follow the original parameters for full parameter training. Additional implementation details are detailed in Appendix 2 and Appendix 6.
@@ 4.2 Component Analysis of Alignment
4.2.1 Visual encoder selection on alignment
To evaluate how visual encoder selection impacts cross-modal alignment and generation quality, we conduct experiments with three encoders: BioMedCLIP [61] (domain-specific cross-modal encoders), CLIP [38] (general-purpose cross-modal encoders), and Med-SAM [33] (domain-specific pure visual encoders). As shown in Table 1 and Appendix 5, all pretrained encoders improve generation performance compared to training without alignment. This improvement arises from the global semantic constraints imposed by external representations, which regularize the autoregressive generation process. Notably, for training from an text-only generative LLM (ARRA), BioMedCLIP and CLIP demonstrate superior performance, as their cross-modal training bridges the gap between text and image modalities. This enables the LLM to learn what to generate (semantics) before how to generate (pixel details). Conversely, when fine-tuning pretrained LLMs with image-generation capabilities (ARRA-Adapt), domain-specific encoders like BioMedCLIP and MedSAM dominate. BioMedCLIP injects medical-specific semantics, while MedSAM provides structural priors (e.g., organ shapes) through its segmentation-focused features.
These results lead to the following key insight: Takeaway 1. When the LLM lacks image capabilities, cross-modal encoders are crucial for semantic grounding. However, for pretrained LLMs with image capabilities, domain-specific encoders are more effective, as they provide the fine-grained features needed for domain-specific adaptation.
Table 1: Impact of alignment with representation extracted from different encoders on MIMIC-CXR, where the gray part represents the ARRA model, and the white part represents the ARRA-Adapt.
Target Rep. FID $\downarrow$ MS-SSIM $\uparrow$ CLIP-Score $\uparrow$
BioMedCLIP [61] 4.15 0.422 0.4576
Med-SAM [33] 4.08 0.398 0.4542
CLIP-L [38] 4.63 0.407 0.4519
w/o align. 5.10 0.401 0.4518
BioMedCLIP [61] 5.30 0.405 0.4532
Med-SAM [33] 6.54 0.384 0.4465
CLIP-L [38] 5.16 0.394 0.4450
w/o align. 7.11 0.383 0.4460
Table 2: Impact of different alignment mechanism selection on MIMIC-CXR. Gray part represents the ARRA model, and the white part represents the ARRA-Adapt.
@@ 4.2.1 Visual encoder selection on alignment
Table 3: Impact of aligning representation to different layers on MIMIC-CXR. Gray part represents the ARRA, and the white part represents the ARRA-Adapt.

### v4

@@ 3.3 Autoregressive Representation Alignment
We align the visual representations extracted from the pretrained foundational model with LLM’s hidden states and investigate impact of different alignment strategies. The goal of alignment is to enable the hidden state of autoregressive transformer to acquire external global representations, providing meaningful guidance for reconstructing image.
Pre-trained Visual Representation Extraction. Let $\mathcal{E}_{F}$ be a pretrained foundation model’s visual encoder and $I$ be a target image. We encode $I$ as a visual representation by $f_{F}=\mathcal{E}_{F}(I)\in\mathbb{R}^{N\times D}$, where $N,D$ denotes the embedding length and dimension of $f_{F}$. $f_{F}$ is aggregated to the global visual representation $f_{GF}$, i.e., $f_{GF}=\text{agg}(f_{F})\in\mathbb{R}^{1\times D}$, where $\text{agg}(\cdot)$ denotes a feature aggregation operation. This aggregation operation fully extracts global information in the features and facilitates alignment with the hidden states of autoregressive model. For CLIP series, inspired by (raghu2021vision), we use the <CLS> token representation from the Transformer-based visual encoder as global visual representation. For SAM series, which lack a <CLS> token, we instead apply average pooling over all patch features for feature aggregation operation.
Hybrid Next Token. We define the next token predicted by LLM sequence in our framework as “HYBRID next token”, denoted as <HYBNEXT>. Unlike a “locally constrained token” in previous autoregressive models that are solely constrained by the LLM codebook, our <HYBNEXT> can fully incorporate external, well-trained global visual representations, making it a “globally and locally constrained token”.
Global Visual Representation Alignment. We obtain the hidden state $f_{L}^{i}$ of <HYBNEXT> from the autoregressive model $\mathcal{M}_{\theta}$. The hidden state $f_{L}^{i}$ is converted to $f_{A}\in\mathbb{R}^{1\times D}$ by a projection layer $\mathcal{A}_{\phi}$ to align with the global visual representation $f_{GF}\in\mathbb{R}^{1\times D}$, i.e., $f_{A}=\mathcal{A}_{\phi}(f_{L}^{i})$, where $\mathcal{A}_{\phi}$ is a two-layer MLP. Representation alignment is achieved through a Global Visual Alignment loss $\mathcal{L}_{\text{GVA}}$, which maximizes the similarity between the projected feature $f_{A}$ and the global visual representation $f_{GF}$:
@@
where sim($\cdot,\cdot$) denotes cosine similarity loss. This alignment enables the <HYBNEXT> token to learn global visual representation, bridging the cross-modal gap and making the token prediction process more reliable.
@@
3.4 Versatile ARRA for Diverse Scenarios
Our ARRA framework is flexible and plug-and-play, supporting different training scenarios and LLM frameworks. Therefore, we provide three representative model variants:
(1) ARRA-Base. It trains an LLM from scratch with random initialization, supporting settings where no pretrained models are available and showcasing ARRA’s ability to learn multimodal alignment from the ground up.
(2) ARRA. It initializes with a pretrained LLM that has strong text generation capabilities, enabling efficient extension to text-to-image tasks with text-generation-only LLMs.
(3) ARRA-Adapt. It builds on pretrained LLMs with both text and image generation capabilities, allowing adaptation to specialized domains such as medical imaging by leveraging special-domain priors.
4 Experimental Analysis and Results
We first perform a comprehensive component analysis of the ARRA framework to investigate how different design choices, such as alignment mechanism, feature aggregation strategy, and external encoder selection, impact alignment performance. Based on analysis, we finalize the ARRA framework and compare the three model variants, ARRA-Base, ARRA, and ARRA-Adapt, with advanced generative models to evaluate their performance and adaptability.
@@ 4.1 Experimental Setup
Figure 3: ARRA-Base improves the generation of LlamaGen. Left are text-conditional image generation results on the LAION-COCO dataset, and right are category-conditional image generation results on the Imagenet dataset.
Datasets. We evaluate our model on both natural and medical image datasets. For natural images, we conduct evaluations on the ImageNet (imagenet) dataset and a 2.4M high-quality subset of LAION-COCO (schuhmann2022laion). For medical imaging, we use the MIMIC-CXR (johnson2019mimic) and DeepEyeNet (Huang2021DeepEyesNet) datasets. Detailed preprocessing is provided in Appendix 3.
Implementation Details. We adopt LlamaGen (sun2024autoregressivemodelbeatsdiffusion), Chameleon 7B (team2024chameleon) and Lumina-mGPT 7B (liu2024lumina) as respective baseline models for the ARRA-Base, ARRA, and ARRA-Adapt model variants, respectively, to verify the effectiveness of our framework.
@@ 4.2 Component Analysis of Alignment
To better understand the core design principles of ARRA and build an effective model, we analyze how different alignment component designs affect the effectiveness of representation alignment in autoregressive generation. We focus on the following key questions:
  • Alignment mechanism: Does token-level alignment (our proposed <HYBNEXT>) outperform fixed-position alignment (<REP>) in preserving global visual constraints during autoregressive generation? (Table 1)
  • Feature aggregation strategy: How do different types of features extracted from the same visual encoder affect the generation performance? (Table 2)
  • Visual encoder selection on alignment: How does external encoder selection (general vs. specialized and cross-modal vs. vision-only) impact generation? (Table 3)
@@
Table 1: Impact of different alignment mechanism selection on MIMIC-CXR. Gray represents ARRA-adapt model, and white represents ARRA model.
Alignment mechanism.
We compare two strategies for integrating visual representations: (1) aligning features to a fixed <REP> token at the start of the generated sequence, and (2) aligning to the hidden state of <HYBNEXT>, a hybrid token interleaved at every generation step. As shown in Table 1, <HYBNEXT> yields superior performance. We argue that <HYBNEXT> allows for comprehensive traversal of every token during training sampling, ensuring that each token is effectively constrained by external global representations. In contrast, <REP> suffers from the “attention sink” (liu2024lumina; xiao2023efficient) effect, where attention to the fixed token decays over long sequences, leading to degraded outputs. This leads to a key insight: Takeaway 1. Aligning visual representations to a hybrid token interleaved at each generation step <HYBNEXT> is more effective than using a fixed token <REP>, as it prevents attention decay and ensures consistent constraint by external representations.
CLS Avgpool FID $\downarrow$ MS-SSIM $\uparrow$ CLIP-Score $\uparrow$
5.30 0.405 0.4532
6.56 0.387 0.4434
5.93 0.385 0.4465
Feature aggregation strategy.
To investigate how different types of features extracted from the visual encoder affect the generation performance, we understand three strategies for aggregating features from a single visual encoder: (1) the [CLS] token representation, (2) average pooling of all image patch representations, and (3) a concatenation of both representations. As shown in Table 2, the [CLS] token representation yields optimal performance. We attribute this superiority to the [CLS] token’s ability to aggregate global visual information through self-attention mechanisms. This ability provides a compact yet comprehensive representation for cross-modal alignment. This finding consistent with (raghu2021vision), which establishes that [CLS] tokens capture global representations while patch tokens focus on local features. This global-local distinction suggests that autoregressive image generation models particularly benefit from external global semantic representations as guidance. These results lead to a key insight: Takeaway 2. The [CLS] token representation in foundation models effectively aggregates global visual information, providing comprehensive guidance for cross-modal alignment.
Target Rep. FID $\downarrow$ MS-SSIM $\uparrow$ CLIP-Score $\uparrow$
BioMedCLIP (zhang2023biomedclip) 4.15 0.422 0.4576
Med-SAM (MedSAM) 4.08 0.398 0.4542
CLIP-L (radford2021CLIP) 4.63 0.407 0.4519
w/o align. 5.10 0.401 0.4518
BioMedCLIP (zhang2023biomedclip) 5.30 0.405 0.4532
Med-SAM (MedSAM) 6.54 0.384 0.4465
CLIP-L (radford2021CLIP) 5.16 0.394 0.4450
w/o align. 7.11 0.383 0.4460
Table 3: Impact of alignment with representation extracted from different encoders on MIMIC-CXR.
Visual encoder selection on alignment.
To evaluate how visual encoder selection impacts cross-modal alignment and generation quality, we conduct experiments with three encoders: BioMedCLIP (zhang2023biomedclip) (domain-specific cross-modal encoders), CLIP (radford2021learning) (general-purpose cross-modal encoders), and Med-SAM (MedSAM) (domain-specific pure visual encoders). As shown in Table 3, all pretrained encoders improve generation performance compared to without alignment. This improvement arises from global constraints imposed by external representations, which regularize the autoregressive generation process. Notably, for training from a text-generation-only LLM (ARRA), BioMedCLIP and CLIP demonstrate superior performance. Their cross-modal training helps bridge the gap between text and image modalities, enabling LLMs to learn what to generate (semantics) before how to generate (pixel details). Conversely, when fine-tuning pretrained LLMs with image-generation capabilities (ARRA-Adapt), domain-specific encoders like BioMedCLIP and MedSAM dominate. BioMedCLIP injects medical-specific semantics, while MedSAM provides structural priors (e.g., organ shapes) through its segmentation-focused features. These results lead to a key insight: Takeaway 3. When the LLM lacks image generation capabilities, cross-modal encoders are crucial for semantic grounding. However, for LLMs with image generation capabilities, domain-specific encoders are more effective, as they provide fine-grained features needed for domain-specific adaptation.
4.3 Main Comparison Results
Figure 4: Visual comparison of different methods with ARRA and ARRA-Adapt on the MIMIC-CXR dataset.
Based on analysis, we finalize ARRA framework with CLIP and BioMedCLIP for alignment on natural and medical datasets, respectively. We use [HYBNEXT] and [CLS] as our alignment strategy. We compare ARRA framework under the three proposed model variants (ARRA-Base, ARRA, ARRA-Adapt) to demonstrate its superiority and versatility.
ImageNet LAION-COCO
Model  [^1] M 343M Model 343M 775M 1.4B 3.1B
FID ↓ IS ↑ FID ↓ IS ↑ (FID, CLIP-Score) (FID, CLIP-Score) (FID, CLIP-Score) (FID, CLIP-Score)
LlamaGen 5.47 193.6 4.33 286.6 LlamaGen (12.78, 0.2470) (11.79, 0.2512) (11.57, 0.2510) (11.88, 0.2605)
ARRA-Base 5.08 201.2 3.61 262.6 ARRA-Base (11.67, 0.2500) (10.81, 0.2571) (10.50, 0.2589) (10.45, 0.2615)
Table 4: Quantitative performance comparison of ARRA-Base and LlamaGen models with different scales on ImageNet datasets and 2.4M subset of LAION-COCO datasets, respectively.
Chest-Xray: MIMIC-CXR Fundus: DeepEyeNet
Method FID $\downarrow$ IS $\uparrow$ MS-SSIM $\uparrow$ CLIP-Score $\uparrow$ FID $\downarrow$ MS-SSIM $\uparrow$ CLIP-Score $\uparrow$
Original SD V2-1(rombach2022high) 71.68 2.365 0.128 0.2333 166.45 0.141 0.2164
DreamBooth SD (ruiz2023dreambooth) 60.40 2.269 0.270 0.3693 80.70 0.352 0.3061
MINIM (wang2025self) 15.62 2.468 0.317 0.4423 59.71 0.333 0.2862
UniXGen (lee2024vision) 30.75 2.437 0.361 0.4128 - - -
LLM-CXR (lee2024llm) 5.88 2.134 0.395 0.4374 - - -
Chameleon (team2024chameleon) 7.11 2.498 0.383 0.4460 38.37 0.341 0.3234
ARRA 5.30 2.587 0.405 0.4532 35.01 0.376 0.3354
ARRA-Adapt 4.15 2.746 0.422 0.4576 34.70 0.392 0.3389
Table 5: Compare different methods with ARRA and ARRA-Adapt on MIMIC-CXR and DeepEyeNet datasets, respectively.
Facilitating training T2I LLMs from scratch: performance comparison of ARRA-Base. We evaluate the generation performance and model scalability of ARRA-Base variant, which is trained from scratch, on large-scale datasets. We first conduct experiments on ImageNet, using LlamaGen (sun2024autoregressivemodelbeatsdiffusion) with  [^1] M and 343M parameters as baselines. We then train models on a 2.4M high-quality subset of LAION-COCO and employ LlamaGen variants of increasing sizes (343M to 3.1B parameters) as baselines. As shown in Table 4 and Fig. 3, ARRA-Base outperforms the baseline model across all scales. The FID decreased by 16.6% and 12.0% on the ImageNet and LAION-COCO datasets, respectively. Meanwhile, as the model size increased, the performance remained steadily improved (FID decreased from 11.67 to 10.45), demonstrating ARRA maintains strong scalability of autoregressive model. These results lead to a key insight: Takeaway 4. ARRA-Base enables efficient training of T2I LLMs from scratch, while preserving strong model scalability.
Boosting pretrained LLMs for T2I generation and domain adaptation: performance of ARRA and ARRA-Adapt. We evaluate ARRA and ARRA-Adapt, initialized with text-generation-only and general image generation pretrained LLMs, respectively, on the MIMIC-CXR and DeepEyeNet datasets. Both variants are benchmarked against state-of-the-art diffusion models (Stable Diffusion (rombach2022high), DreamBooth (ruiz2023dreambooth), MINIM (wang2025self)) and autoregressive models (LLM-CXR (lee2024llm), UniXGen (lee2024vision), Chameleon (team2024chameleon)). As shown in Table 5, our model achieves superior performance in both visual quality and semantic alignment. Compared to the chameleon, ARRA achieves a 25.5% and 8.8% reduction in FID on the MIMIC-CXR and DeepEyeNet datasets, respectively. On the MIMIC-CXR dataset, ARRA surpasses the best-performing baseline, LLM-CXR, reducing the FID from 5.88 to 5.30 and increasing the CLIP-Score from 0.4374 to 0.4532. The ARRA-Adapt variant delivers even stronger results, achieving an FID of 4.15 and a CLIP-Score of 0.4576. In addition, as shown in Fig. 4, our models demonstrate superior alignment with fine-grained clinical details, such as lesion location and severity. These results lead to a key insight: Takeaway 5. ARRA enables a more effective transformation of text-generated-only LLMs into T2I generators, while ARRA-Adapt substantially improves domain adaptation and aligns general image-generation LLMs with specialized fields more effectively, both outperforming baseline approaches.
5 Conclusion
We propose Autoregressive Representation Alignment (ARRA), a framework that enhances autoregressive image generation by injecting external visual representations during training. This approach enriches global semantic understanding while maintaining the model’s original autoregressive paradigm during generation. Experiments on natural and medical image generation tasks demonstrate ARRA’s versatility, offering a cost-effective framework for training autoregressive text-to-image generation models. Our work bridges the gap between multimodal domains and provides novel insights into modeling unified multimodal generation.
6 Acknowledgments
This work was supported by the National Natural Science Foundation of China (62306253, 61873259), the Early Career Fund (27207025), the National Key Research and Development Program of China under Grant 2024YFB4707700, the National Natural Science Foundation of China under Grant U23A20343 and the Guangdong Natural Science Fund-General Program (2024A1515010233).
Figure 1: More visualizations on the chest X-ray generation task across the MIMIC-CXR datasets.
1 Overview
Our supplementary materials include the following sections:
  • Discussion with related work (Section 2).
  • Additional implementation details about ARRA and other comparative methods in MIMIC-CXR and DeepEyeNet (Section 3);
  • Implementation details of component analysis of alignment. (Section 4);
  • Ablation study, including generalizability at different resolutions, choices of projection layers, regularization coefficients $\lambda$, training objectives and alignment depth (Section 5);
  • More visual results on the MIMIC-CXR dataset (Section 6);
  • Implementation details and more results on the LAION-COCO and ImageNet datasets(Section 7);
2 Discussion with related work
Discussion with REPA (yu2025repa): While both works leverage feature alignment concepts, ARRA presents fundamentally distinct contributions from REPA:
Different motivation: REPA resolves challenges in learning high-quality internal representations for diffusion models by injecting representations from external clean images. In sharp contrast, ARRA addresses a fundamental limitation of autoregressive LLMs in text-to-image generation: the inherent inability to capture global coherence. Unlike prior works requiring complex architectural redesigns to impose global constraints, ARRA demonstrates that such global coherence naturally emerges through our alignment-driven training paradigm, achieved by aligning LLM hidden states with visual representations from an external visual encoder.
[…]
3 Additional implementation details in MIMIC-CXR and DeepEyeNet
Datasets. For the MIMIC-CXR, we employ three distinct radiographic views: posteroanterior (PA), anteroposterior (AP), and lateral (LATENT). Input prompts are generated from the “impression” section of diagnostic reports, formatted as: “{view} view chest X-ray image, {impression}”. We select 221,238 pairs for training and 1,000 pairs for testing. DeepEyeNet (Huang2021DeepEyesNet) is a comprehensive fundus dataset, containing more than 15k high-quality “text-image” pairs. In our experiments, we derive prompts from the “clinical-description” section and selectively choose 7,190 training pairs and 1,089 testing pairs. All images are preprocessed through center cropping to 512×512 pixels.
Evaluation Metrics. We employ three quantitative metrics to evaluate the proposed method: Fréchet Inception Distance (FID), Multi-Scale Structural Similarity (MS-SSIM), and CLIP-Score. For X-ray image generation, FID is computed using features extracted from a CXR-pre-trained DenseNet-121 (XRV) (cohen2022torchxrayvision), while for fundus images, we utilize features from an ImageNet-pre-trained Inception V3 (heusel2017gans). The CLIP-Score metric is derived from the cosine similarity between image-text feature pairs encoded through BioMedCLIP (zhang2023biomedclip), a large-scale medical vision-language model.
ARRA and ARRA-Adapt. All experiments on medical images are implemented using PyTorch on 4 NVIDIA A6000 Ada GPUs. We employ a batch size of 6 with a learning rate of 2e-5, optimized by AdamW (weight decay=0.05, $\beta_{1}$=0.9, $\beta_{2}$=0.95). To stabilize training, we apply z-loss regularization (team2024chameleon) with a weight of 1e-5. All data are pre-tokenized before training to increase throughput. The VQ tokenizer operates with a downsampling rate of 16, resulting in a 1024-token representation for each image.
Original SD V2-1. (rombach2022high) Stable-diffusion-2-1 is a diffusion model trained from natural images, and we obtain its official pre-training weights, which are used directly to generate images without training, to demonstrate its zero-shot generation capability.
DreamBooth SD. (ruiz2023dreambooth) We adapt Dreambooth to fine-tune the stable-diffusion-xl-base model, choosing “X-rays” and “fundus” as class identifiers [class noun] on the MIMIC-CXR and DeepEyesNet datasets, respectively, and supervised the fine-tuning using the full data.
MINIM. (wang2025self) MINIM is a medical image generation model based on stable-diffusion-v1-4 for full fine-tuning. We use its official implementation for training and generation.
UniXGen. (lee2024vision) The UniXGen model is a unified model for X-ray image understanding and generation, fully trained on the MIMIC-CXR dataset. We use its official implementation and obtain its pre-training weights for generation.
LLM-CXR. (lee2024llm) The LLM-CXR model is a unified model for X-ray image understanding and generation using dolly-v2-3b (DatabricksBlog2023DollyV2) as a baseline for full parameter fine-tuning training. We implement training and inference on the MIMIC-CXR dataset using its official implementation.
Chameleon. (team2024chameleon) Chameleon is a multimodal large language model trained on the natural image datasets. Its open-source weights are only for visual language understanding. We follow its official implementation and use 7B pre-training weights to fine-tune full parameters for training on the x-ray and fundus image dataset.
4 Implementation of component analysis
For the Alignment mechanism, we align the representations extracted from BioMedCLIP.
For the Feature aggregation strategy, we align the representations extracted from the BioMedCLIP model to the hidden state of <HYBNEXT>.
For the Visual encoder selection on alignment, we align the representations to the hidden state of <HYBNEXT>.
5 More Ablation study
Generalizability at different resolutions.
We validate the generalization capability of the ARRA framework at resolutions of 256×256 and 512×512, as shown in Table 1. Experimental results demonstrate that ARRA can stably improve the generation performance of baseline models at different resolutions, demonstrating its resolution generalization ability.
Table 1: Generalizability at different resolutions.
Resolution Model FID $\downarrow$ MS-SSIM $\uparrow$ CLIP-Score $\uparrow$
256×256 Chameleon 9.42 0.369 0.4405
256×256 ARRA 6.11 0.388 0.4436
512×512 Chameleon 7.11 0.383 0.4460
512×512 ARRA 5.30 0.405 0.4532
Table 2: Compare performance of different alignment heads.
MLP Maxpool FID $\downarrow$ MS-SSIM $\uparrow$ CLIP-Score $\uparrow$
5.30 0.405 0.4532
5.37 0.386 0.4375
Projection layer.
We evaluate the performance of different projection layers, including a two-layer MLP and a direct pooling layer. As shown in Table 2, our experiments reveal that using a two-layer MLP outperforms the pooling layer, even when the projection layer is discarded during testing. We hypothesize that, compared to the non-trainable pooling layer, employing a trainable MLP as a ”soft link” during training allows the model to better learn information-rich and precise representations while adaptively filtering out irrelevant information.
Effect of $\lambda$.
We also examine the effect of the regularization coefficient $\lambda$ by training models with different coefficients from 0.5 to 2 and comparing the performance. As shown in Table 3, the optimal value is reached when the regularization coefficient $\lambda$ = 1.
Table 3: Ablation study for alignment loss coefficient $\lambda$.
$\lambda$ FID $\downarrow$ MS-SSIM $\uparrow$ CLIP-Score $\uparrow$
0.5 6.22 0.389 0.4468
0.8 5.38 0.399 0.4460
1 5.30 0.405 0.4532
1.5 5.87 0.394 0.4477
2 6.12 0.390 0.4479
Training objective.
We compare two alignment objectives: MSE loss and cosine similarity loss. As shown in Table 4, both achieve similar FID and MS-SSIM scores, but cosine similarity yields better semantic accuracy (higher CLIP-Score). This advantage stems from its focus on directional similarity rather than absolute distances, making it more robust for cross-modal alignment tasks.
Table 4: Impact of different alignment objectives.
Objective FID $\downarrow$ MS-SSIM $\uparrow$ CLIP-Score $\uparrow$
Cos. sim. 5.30 0.405 0.4532
MSE 5.32 0.405 0.4454
Alignment depth.
These results lead to a key insight: Takeaway 4. Performing representation alignment at earlier LLM layers provides sufficient global semantic guidance while allowing deeper layers to focus on capturing high-frequency details.
Table 5: Impact of aligning representation to different layers on MIMIC-CXR. Gray part represents the ARRA-adapt, and the white part represents the ARRA.